OpenStudy (anonymous):

evlutate :( (1-u)+1)^2

OpenStudy (anonymous):

this is part of an intergral i am doing

OpenStudy (anonymous):

intergral: (x+1)^2(1-x)^5

OpenStudy (anonymous):

isn't this \[(2-u)^2\]?

OpenStudy (anonymous):

i choose u as being (1-x) and then reworte x in terms of u

OpenStudy (anonymous):

but x=1-u

OpenStudy (anonymous):

then i subs it in to (x+1)^2

OpenStudy (aravindg):

satellite help meeee

OpenStudy (anonymous):

i think that is just swapping apples for oranges

OpenStudy (anonymous):

so, then how would i go about solving this, i got no clue

OpenStudy (anonymous):

do i just have to expand

OpenStudy (anonymous):

think best bet is to grind it til you find it multiply all that crap out, combine like terms, integrate term by term, cheat and uses a machine

OpenStudy (anonymous):

but, there has to be another way i am not seeing, cause this is a problem from the solving integrals by sub chp

OpenStudy (anonymous):

after i expand each term: (x+1)^2 and (1-x)^5 i have to multiply them together right?

OpenStudy (anonymous):

\[u=x+1\] \[1-x=1-(u+1)=2-u\] and now you have \[u^2(2-u)^5 which buys you nothing

OpenStudy (anonymous):

let me think before i speak

OpenStudy (anonymous):

zarkon?

OpenStudy (aravindg):

sate

OpenStudy (zarkon):

how about u=1-x that would make multiplying it out a little easier

OpenStudy (zarkon):

I don't see a really nice way at first glance

OpenStudy (anonymous):

man , but in the end i will still have to sub back in that 1-x and multiply it out

OpenStudy (zarkon):

I would sub back 1-x but I wouldn't multiply it all out

OpenStudy (anonymous):

so i would get: (x^2+2x+1)(u)^5

OpenStudy (anonymous):

oh yes zarkon is as usual right. it is just a question of figuring out the cheapest way to do the multiplication

OpenStudy (anonymous):

\[u=1-x\] \[x+1=2-u\] and then \[(2-u)^2u^5=(4-4u+u^2)u^5=4u^5-4u^6+u^7\] but this doesn't really help either does it because you still have to either integrate and then expand, or expand and then integrate. sorry

OpenStudy (anonymous):

how did you get that: x+1=2-u

OpenStudy (zarkon):

what is wrong with integrating ...substituting back 1-x and then leaving as it is ...it will be the anti derivative.

OpenStudy (anonymous):

sure why not. everything will just be in terms of 1- x instead of x

OpenStudy (anonymous):

\[u=1-x\] \[x=1-u\] \[x+1=1-u+1=2-u\] is the answer to your question

OpenStudy (anonymous):

got you homie

OpenStudy (anonymous):

let me get down to work